# RankCLIP: Ranking-Consistent Language-Image Pretraining

### v2

@@ Abstract
RankCLIP: Ranking-Consistent Language-Image Pretraining
Yiming Zhang††thanks: These authors contribute equally. Correspondence to Zenghui Ding <dingzenghui@iim.ac.cn>. Zhuokai Zhao Zhaorun Chen2 Zhili Feng3 Zenghui Ding1 Yining Sun1 1HFIPS, Chinese Academy of Sciences 2University of Chicago 3Carnegie Mellon University 4University of Science and Technology of China
@@ 1 Introduction
In the realm of computer vision (CV) (Voulodimos et al., 2018), natural language processing (NLP) (Chowdhary and Chowdhary, 2020), and multimodal deep learning (Jabeen et al., 2023; Zhao et al., 2023), the alignment between visual and textual modalities (Singh et al., 2022; Chen et al., 2024a) has emerged as a cornerstone for downstream applications, ranging from image captioning (Ghandi et al., 2023) to zero-shot classification (Pourpanah et al., 2022). Contrastive Language-Image Pretraining (CLIP) (Radford et al., 2021) marks a significant advancement in this field, demonstrating incredible performance from training on large amounts of text-image pairs to create self-supervised models that understand (Hendrycks et al., 2021a, b; Chen et al., 2024b) and generate (Ramesh et al., 2021; Crowson et al., 2022) descriptions of visual contents.
@@
Following the success of this contrastive learning paradigm, many recent works have been developed upon the original CLIP. More specifically, these enhancements focus on optimizing data efficiency through intrinsic supervision (Li et al., 2021), as well as improving downstream performance via cross-modal late interaction (Yao et al., 2021), hierarchical feature alignment (Gao et al., 2022), geometric consistency regularization (Goel et al., 2022), additional learning (Mu et al., 2022), adaptive loss (Yang et al., 2023), hierarchy-aware attentions (Geng et al., 2023), and softer cross-modal alignment (Gao et al., 2024).
@@
Despite the improvements, these methods often have reliance on strict pairwise, cross-modal, and one-to-one mappings between images and texts, overlooking the actual many-to-many relationships that exist both cross-modal and in-modal in real-world data (Chun, 2023). For example, as shown in Fig. 1, while pretrained models like CLIP can correctly classify dog, cat and airplane, they do not necessarily learn that dog and cat are more close to each other than dog and airplane, in terms of both in-modal (dog text is more similar to cat text than to airplane text) and cross-modal (dog text is more matched to cat image than to airplane image) similarities. Because it is rooted from the current contrastive loss that only the correct pairs are optimized while the rest of the unmatched pairs are treated the same, resulting in a large amount of information not used and unknown to the model during and after the training process.
@@
Recognizing the complex many-to-many relationships as well as the rich information contained within both in-modal and cross-modal data, we introduce Ranking-Consistent Language-Image Pretraining, (RankCLIP), which employs ranking consistency to learn and optimize similarity levels both between (cross-modal) and within (in-modal) the text-image pairs. The concept of ranking consistency stems from the simple observations that similar texts often correlate with similar images, as seen with the dog, cat and airplane example in Fig. 1. It effectively captures secondary similarity relationships among unmatched pairs, enabling the model to learn more efficiently for free compared to relying solely on matched pairs. Ranking consistency is conveniently modeled as an additional loss term to the traditional contrastive loss, requiring no extra external modules. It acts as a plug-and-play improvement for many existing methods, including those focusing on data-efficiency (Li et al., 2021), potentially boosting performance in both efficiency and effectiveness.
@@
The main contributions of this paper are: (1) RankCLIP, a novel contrastive pretraining method that uses ranking consistency to exploit the many-to-many relationships within data, thereby enhancing performance in downstream tasks such as zero-shot classification and retrieval accuracy; and (2) through comprehensive experiments conducted on multiple datasets, we demonstrate the superior effectiveness of RankCLIP in improving pretraining model performance without requiring any additional data or computational resources.
@@ 2 Related Work
Vision-language pretraining has witnessed significant advancements over the past years (Chen et al., 2023; Du et al., 2022; Long et al., 2022). Models such as CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021) and FLAVA (Singh et al., 2022) have pioneered the contrastive learning paradigm applied with text-image pairs, showcasing remarkable performance and robustness in downstream tasks. Many follow-up works, mostly built upon CLIP, have been proposed since then. Li et al. (2021) introduced DeCLIP, improving zero-shot performance through intrinsic supervision. FILIP (Yao et al., 2021) advances CLIP’s alignment between image patches and text with a cross-modal interaction mechanism. Gao et al. (2022) developed PyramidCLIP, using hierarchical feature alignment to boost model efficiency and performance. Additionally, SLIP (Mu et al., 2022) merges self-supervised learning with CLIP pre-training for improved visual representation and accuracy. Goel et al. (2022) introduced CyCLIP, augmenting CLIP with geometric consistency regularizers to enhance robustness and performance under varied conditions.
Recently, Yang et al. (2023) introduced ALIP, an adaptive pre-training model that enhances language-image alignment using raw text and synthetic captions with dynamic adjustments. HiCLIP (Geng et al., 2023) refines CLIP by adding hierarchy-aware attentions to uncover semantic hierarchies in images and texts. EqSim (Wang et al., 2023) incorporates equivariance loss into vision-language models, significantly improving sensitivity to semantic changes in image-text pairs. Additionally, SoftCLIP (Gao et al., 2024) softens CLIP’s one-to-one constraint, enabling more flexible cross-modal alignment through fine-grained adjustments.
@@ 3 RankCLIP
RankCLIP efficiently leverages the many-to-many relationships in real-world data by focusing on both matched and unmatched pairs. As shown in Fig. 2, it not only identifies if an image-text pair matches but also assesses their relative semantic similarities to other images and texts of both modalities in the dataset through self-supervised ranking consistency. Uniquely, RankCLIP employs a list-wise loss for training batches, distinguishing it from other methods that solely rely on pair-wise relationships, as discussed in §2.
@@ 3.1 Ranking Model Formulation
RankCLIP leverages the Plackett-Luce (PL) ranking model Plackett (1975); Luce (2005); Guiver and Snelson (2009) to estimate the probability distribution over rankings for every image-text pair , so that the consistency in their relative ordering with respect to a reference ranking can be measured. Specifically, for a given data pair, whether it is in-modal (image-image, text-text), or cross-modal (image-text), we calculate its in- or cross-modal cosine similarity to serve as the score when measuring the alignment of its ranking with respect to another reference ranking .
Following Plackett (1975), we first sort the reference ranking in a descending order to construct the optimal ranking , and assume that the ego ranking is sampled from . Thus the probability that item with score is ranked in the ego ranking from a set of items is the score of divided by the sum of scores for the items that have not been placed yet:
where denotes the set of items ranked before . In addition, we propose a decaying factor to scale the loss, so that the top-ranked items can obtain higher weights: Consequently, the probability of the entire ranking is the product of all individual placement probabilities:
@@ 3.4 RankCLIP Loss
Combining both cross-modal and in-modal consistency with the traditional contrastive loss (more details in Appendix A), the complete rank loss is thus formulated as:
which is also depicted in Fig. 2. By supplementing the pairwise contrastive loss with cross-modal and in-modality ranking consistency loss, RankCLIP systematically organizes embeddings to fully exploit both global and fine-grained secondary unmatched relationships, which enhances the learning of more informative and accurate representations, better supporting downstream multi-modal tasks.
| CIFAR-10 | CIFAR-100 | ImageNet1K | |||||||
| Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | |
| CLIP | 36.35% | 70.28% | 85.02% | 12.22% | 24.93% | 33.56% | 12.08% | 21.86% | 27.48% |
| ALIP | 35.71% | 72.39% | 88.77% | 13.67% | 27.10% | 34.76% | 15.62% | 26.90% | 32.50% |
| RankCLIP | 37.03% (+1.87%) | 67.67% (-3.71%) | 83.09% (-2.27%) | 13.98% (+14.40%) | 27.70% (+11.11%) | 36.17% (+7.78%) | 17.02% (+40.89%) | 28.44% (+30.10%) | 33.99% (+23.69%) |
| ImageNetV2 | ImageNetSketch | ImageNet-A | ImageNet-R | |||||||||
| CLIP | 12.11% | 22.66% | 28.57% | 3.20% | 7.00% | 9.83% | 3.16% | 8.81% | 13.04% | 11.34% | 21.38% | 27.10% |
| ALIP | 15.62% | 27.34% | 32.82% | 5.10% | 10.37% | 14.01% | 3.53% | 9.14% | 13.61% | 14.25% | 25.74% | 32.43% |
| RankCLIP | 17.03% (+40.63%) | 28.60% (+26.21%) | 34.18% (+19.64%) | 5.82% (+81.88%) | 11.35% (+62.14%) | 14.87% (+51.27%) | 3.82% (+20.89%) | 9.16% (+3.97%) | 13.77% (+5.60%) | 15.74% (+38.80%) | 27.51% (+28.67%) | 34.36% (+26.79%) |
@@ 4.1 Experimental Setup
Baselines.
The most direct baseline to RankCLIP is the original CLIP (Radford et al., 2021). In addition, to further demonstrate the superior performance of RankCLIP, we include ALIP (Yang et al., 2023), which leverages synthetic captions to enhance vision-language representation learning. More specifically, it employs a unique architecture that dynamically adjusts sample and pair weights to mitigate the impact of noisy or irrelevant data, which is quite orthogonal to our approach. The training procedures and parameters of all models are detailed in Appendix B.
Pretraining dataset.
Both baseline models, CLIP (Radford et al., 2021), ALIP (Yang et al., 2023) and the proposed RankCLIP are pretrained on the Conceptual Captions 3M (CC3M) dataset (Sharma et al., 2018), which contains around 3.3 million text-image pairs. Despite being much smaller than CLIP’s initial dataset (Ilharco et al., 2021), CC3M effectively supports the development of pretrained models with strong zero-shot capabilities and is widely used in existing language-image pretraining research (Carlini and Terzis, 2021; Li et al., 2021; Tejankar et al., 2021; Mu et al., 2022; Goel et al., 2022).
4.2 Zero-shot Classification
Zero-shot capability is one of the most significant improvements that CLIP achieves. Thus in this section, we first evaluate the zero-shot classification performance of CLIP, ALIP and the proposed RankCLIP. Following (Goel et al., 2022), we conduct our experiments on CIFAR-10 (Krizhevsky et al., 2009), CIFAR-100 (Krizhevsky et al., 2009), and ImageNet1K (Deng et al., 2009; Russakovsky et al., 2015) dataset.
As shown in Table 1, RankCLIP consistently outperforms CLIP across CIFAR-10, CIFAR-100, and ImageNet1K datasets. Relative to CLIP, RankCLIP shows average improvements of +18.95%, +12.5%, and +9.73% in top-1, top-3 and top-5 metrics, respectively. Particularly on the more challenging ImageNet1K dataset, RankCLIP improves relative top-1 accuracy by +40.89% over CLIP, highlighting the effectiveness of the proposed ranking consistency in enhancing language-image alignment and understanding with the same amount of training data. The two cases where RankCLIP does not excel are the top-3 and top-5 accuracy on CIFAR-10. However, this is likely because CIFAR-10 with top-3 and top-5 metrics is much simpler, reducing the demand for a deeper model understanding.
Additionally, we observe that RankCLIP consistently outperforms ALIP, suggesting that our ranking consistency more effectively enhances text-image representations and alignments compared to the synthetic captions proposed in ALIP. Another trend we observe is that RankCLIP shows the most significant improvement in top-1 accuracy compared to top-3 and top-5. Considering the real-world emphasis on the topmost model output, RankCLIP is likely to offer considerable advantages in practical applications.
@@ 4.2 Zero-shot Classification
Average | |
| CLIP | 72.40% | 48.43% | 49.89% | 26.10% | 48.59% | 65.20% | 77.49% | 49.74% | 53.71% | 83.59% | 44.80% | 56.37% |
| ALIP | 73.87% | 51.00% | 58.09% | 27.72% | 49.74% | 60.34% | 73.14% | 59.36% | 53.98% | 87.94% | 38.07% | 57.56% |
| RankCLIP | 72.54% (+0.20%) | 49.16% (+1.50%) | 53.24% (+6.71%) | 24.99% (-4.25%) | 47.11% (-3.05%) | 63.37% (-2.81%) | 86.40% (+11.50%) | 54.10% (+8.77%) | 54.09% (+3.00%) | 86.10% (+0.71%) | 43.30% (-3.35%) | 57.67% (+2.31%) |
| Flickr30K | MSCOCO | |||||||||||
| Text Retrieval | Image Retrieval | Text Retrieval | Image Retrieval | |||||||||
| R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | R@1 | R5 | R@10 | |
| CLIP | 84.00% | 88.70% | 91.0% | 8.70% | 16.90% | 21.20% | 82.06% | 85.24% | 87.82% | 5.04% | 12.98% | 18.32% |
| ALIP | 84.40% | 90.00% | 92.50% | 9.40% | 17.60% | 21.30% | 82.56% | 86.04% | 88.26% | 6.08% | 13.96% | 19.38% |
| RankCLIP | 84.10% (+0.12%) | 89.40% (+0.79%) | 91.90% (+0.99%) | 8.10% (-6.90%) | 16.40% (-2.96%) | 21.70% (+2.36%) | 82.90% (+1.02%) | 85.68% (+0.52%) | 88.00% (+0.20%) | 5.60% (+11.11%) | 13.20% (+1.69%) | 18.02% (-1.64%) |
@@ 4.3 Robustness to Distribution Shifts
To evaluate the robustness of RankCLIP under distribution shifts, we test it alongside CLIP and ALIP across four ImageNet variants, including ImageNetV2 (Recht et al., 2019), ImageNetSketch (Wang et al., 2019), ImageNet-A (Hendrycks et al., 2021b), and ImageNet-R (Hendrycks et al., 2021a), which are designed to assess resilience to different distribution shifts.
As shown in Table 2, RankCLIP outperforms both CLIP and ALIP consistently. Notably, relative to CLIP, RankCLIP’s accuracy improvements in shifted conditions are +45.55% (top-1), +30.24% (top-3), and +25.83% (top-5), surpassing its performance in standard settings (Table 1) of +40.89% (top-1), +30.1% (top-3), and +23.69% (top-5), indicating the even more superior performance in robustness under distribution shifts.

### v3

@@ 1 Introduction
In the realm of computer vision (CV) [55], natural language processing (NLP) [7], and multimodal deep learning [25, 60, 4], the alignment between visual and textual modalities [49, 5] has emerged as a cornerstone for downstream applications, ranging from image captioning [18] to zero-shot classification [42].
@@
Contrastive Language-Image Pretraining (CLIP) [43] marks a significant advancement in this field, demonstrating incredible performance from training on large amounts of text-image pairs to create self-supervised models that understand [22, 23, 6] and generate [44, 11] descriptions of visual contents. Following the success of this contrastive learning paradigm, many recent works have been developed upon the original CLIP. More specifically, these enhancements focus on optimizing data efficiency through intrinsic supervision [28], as well as improving downstream performance via cross-modal late interaction [59], hierarchical feature alignment [15], geometric consistency regularization [19], additional learning [36], adaptive loss [58], hierarchy-aware attentions [17], and softer cross-modal alignment [16].
@@
Despite the improvements, these methods often have reliance on strict pairwise, cross-modal, and one-to-one mappings between images and texts, overlooking the actual many-to-many relationships that exist both cross-modal and in-modal in real-world data [8]. For example, as shown in Fig. 1, while pretrained models like CLIP can correctly classify dog, cat and airplane, they do not necessarily learn that dog and cat are more close to each other than dog and airplane, in terms of both in-modal (dog text is more similar to cat text than to airplane text) and cross-modal (dog text is more matched to cat image than to airplane image) similarities. Since it is rooted from the current contrastive loss that only the correct pairs are optimized while the rest of the unmatched pairs are treated the same, a large amount of information not used and unknown to the model during and after the training process.
@@
Recognizing the complex many-to-many relationships as well as the rich information contained within both in-modal and cross-modal data, we introduce Ranking-Consistent Language-Image Pretraining, (RankCLIP), which employs ranking consistency to learn and optimize similarity levels both between (cross-modal) and within (in-modal) the text-image pairs. The concept of ranking consistency stems from the simple observations that similar texts often correlate with similar images, as seen with the dog, cat and airplane example in Fig. 1. It effectively captures secondary similarity relationships among unmatched pairs, enabling the model to learn more efficiently for free compared to relying solely on matched pairs. Ranking consistency is conveniently modeled as an additional loss term to the traditional contrastive loss, requiring no extra external modules. It acts as a plug-and-play improvement for many existing methods, including those focusing on data-efficiency [28], potentially boosting performance in both efficiency and effectiveness.
@@
The main contributions of this paper are: 1) RankCLIP, a novel contrastive pretraining method that uses ranking consistency to exploit the many-to-many relationships within data, thereby enhancing performance in downstream tasks such as zero-shot classification and retrieval accuracy; and 2) through comprehensive experiments conducted on multiple datasets, we demonstrate the superior effectiveness of RankCLIP in improving pretraining model performance without requiring any additional data or computational resources.
@@ 2 Related Work
Vision-language pretraining has witnessed significant advancements over the past years [3, 14, 31]. Models such as CLIP [43], ALIGN [26] and FLAVA [49] have pioneered the contrastive learning paradigm applied with text-image pairs, showcasing remarkable performance and robustness in downstream tasks. Many follow-up works, mostly built upon CLIP, have been proposed since then. Li et al. [28] introduced DeCLIP, improving zero-shot performance through intrinsic supervision. FILIP [59] advances CLIP’s alignment between image patches and text with a cross-modal interaction mechanism. Gao et al. [15] developed PyramidCLIP, using hierarchical feature alignment to boost model efficiency and performance. Additionally, SLIP [36] merges self-supervised learning with CLIP pre-training for improved visual representation and accuracy. Goel et al. [19] introduced CyCLIP, augmenting CLIP with geometric consistency regularizers to enhance robustness and performance under varied conditions.
Recently, Yang et al. [58] introduced ALIP, an adaptive pre-training model that enhances language-image alignment using raw text and synthetic captions with dynamic adjustments. HiCLIP [17] refines CLIP by adding hierarchy-aware attentions to uncover semantic hierarchies in images and texts. EqSim [57] incorporates equivariance loss into vision-language models, significantly improving sensitivity to semantic changes in image-text pairs. Additionally, SoftCLIP [16] softens CLIP’s one-to-one constraint, enabling more flexible cross-modal alignment through fine-grained adjustments.
@@ 3 RankCLIP
RankCLIP efficiently leverages the many-to-many relationships in real-world data by focusing on both matched and unmatched pairs. As in Fig. 2, it not only identifies if an image-text pair matches but also assesses their relative semantic similarities to other images and texts of both modalities in the dataset through self-supervised ranking consistency. Uniquely, RankCLIP employs a list-wise loss for training batches, distinguishing it from other methods that solely rely on pair-wise relationships, as discussed in §2.
@@ 3.1 Ranking Model Formulation
RankCLIP leverages the Plackett-Luce (PL) ranking model [41, 32, 20] to estimate the probability distribution over rankings for every image-text pair , so that the consistency in their relative ordering with respect to a reference ranking can be measured. Specifically, for a given data pair, whether it is in-modal (image-image, text-text), or cross-modal (image-text), we calculate its in- or cross-modal cosine similarity to serve as the score when measuring the alignment of its ranking with respect to another reference ranking .
Following [41], we first sort the reference ranking in a descending order to construct the optimal ranking , and assume that the ego ranking is sampled from . The probability that item with score is ranked in the ego ranking from a set of items is the score of divided by the sum of scores for the items that have not been placed yet:
where denotes the set of items ranked before . Consequently, the probability of the entire ranking is the product of all individual placement probabilities:
@@ 3.2 Cross-modal Consistency Ranking
ImageNet1K | MSCOCO | ||||||||
Image-to-Text Retrieval | Text-to-Image Retrieval | ||||||||
Top-1 | Top-3 | Top-5 | Recall@1 | Recall@5 | Recall@10 | Recall@1 | Recall@5 | Recall@10 | |
CLIP [43] | 9.06% | 16.94% | 21.63% | 6.68% | 18.36% | 26.94% | 3.70% | 9.74% | 14.04% |
CyCLIP [19] | 9.40% | 17.32% | 21.72% | 6.50% | 19.34% | 29.14% | 3.72% | 11.16% | 16.06% |
ALIP [58] | 9.71% | 18.31% | 23.07% | 6.04% | 18.04% | 26.92% | 3.70% | 10.22% | 14.38% |
RankCLIP | 10.16% | 19.57% | 24.01% | 7.18% | 19.46% | 28.48% | 3.74% | 10.28% | 14.18% |
@@ 3.4 RankCLIP Loss
Combining both cross-modal and in-modal consistency with the traditional contrastive loss (more details in Appendix B), the complete rank loss is thus formulated as:
which is also depicted in Fig. 2. By supplementing the pairwise contrastive loss with cross-modal and in-modality ranking consistency loss, RankCLIP systematically organizes embeddings to fully exploit both global and fine-grained unmatched relationships, which enhances the learning of more informative and accurate representations, better supporting downstream multi-modal tasks.
3.5 Training Recipe on Selecting and
In the early stage of pre-training, rank consistency is highly unstable due to random initialization. Overemphasizing ranking consistency at this stage can impede the optimization of the embedding space. To address this, we gradually increase the weights and of the ranking loss as training progresses. Specifically, we have:
where and denote the current training epoch and total number of epoch, respectively. The full RankCLIP framework is outlined in Algorithm 1.
@@ 4.1 Experimental Setup
Baselines. The most direct baseline to RankCLIP is the original CLIP [43]. To further demonstrate the superior performance of RankCLIP, we include CyCLIP [19], which introduces cyclic consistency constraints to enforce more robust alignment between visual and textual representations, improving generalization and semantic coherence. We also include ALIP [58], which leverages synthetic captions to enhance vision-language representation learning. More specifically, it employs a unique architecture that dynamically adjusts sample and pair weights to mitigate the impact of noisy or irrelevant data, making its approach complementary to ours. The training procedures and parameters for all models are detailed in Appendix A.
Data. All approaches are pretrained on the Conceptual Captions 3M (CC3M) dataset [48], which contains approximately 3.3 million text-image pairs. Although significantly smaller than CLIP’s original dataset [24], CC3M remains a standard benchmark for vision-language pretraining, enabling strong zero-shot performance [2, 28, 53, 36, 19]. As discussed in §5.2, we also train CLIP and RankCLIP on a larger 15M-image subset of YFCC100M [54] (YFCC15M) to study the effect of scaling up the dataset size.
ImageNetV2-Matched | ImageNetV2-Threshold | ImageNetV2-Top | ImageNet-R | |||||||||
CLIP [43] | 7.53% | 14.99% | 19.61% | 8.89% | 17.22% | 21.86% | 10.76% | 19.80% | 24.87% | 9.36% | 10.56% | 19.76% |
CyCLIP [19] | 7.68% | 15.07% | 19.11% | 9.10% | 17.42% | 21.94% | 11.20% | 20.18% | 25.34% | 9.23% | 16.72% | 21.64% |
ALIP [58] | 7.82% | 15.56% | 19.81% | 9.65% | 18.31% | 22.85% | 11.43% | 20.88% | 26.10% | 10.92% | 20.27% | 26.24% |
RankCLIP | 9.01% | 16.95% | 21.12% | 10.32% | 19.31% | 24.13% | 12.31% | 22.11% | 27.17% | 11.34% | 20.88% | 26.94% |
@@
Average |
CLIP [43] | 77.6% | 56.2% | 43.2% | 22.6% | 39.7% | 60.0% | 40.4% | 51.0% | 79.0% | 50.5% | 52.0% |
CyCLIP [19] | 76.8% | 54.3% | 45.8% | 19.2% | 37.5% | 58.6% | 44.2% | 51.5% | 82.3% | 41.3% | 51.2% |
ALIP [58] | 71.1% | 49.1% | 47.1% | 17.4% | 36.1% | 51.5% | 41.9% | 53.3% | 81.0% | 38.3% | 48.7% |
RankCLIP | 78.4% | 56.6% | 42.4% | 23.4% | 40.2% | 60.6% | 40.6% | 53.4% | 79.6% | 47.7% | 52.3% |
4.2 Zero-shot Classification
We evaluate the zero-shot classification performance of CLIP [43], CyCLIP [19], ALIP [58], and RankCLIP on ImageNet1K [12, 46]. As shown in Table 1, RankCLIP consistently outperforms CLIP, which highlight the effectiveness of ranking consistency in enhancing language-image alignment with the same training data. Compared to CyCLIP [19], which enforces cyclic consistency to improve semantic coherence, RankCLIP achieves higher accuracy across all metrics, suggesting that ranking consistency provides a more direct and effective regularization for representation learning. Additionally, RankCLIP surpasses ALIP [58], indicating that ranking consistency is a stronger alternative to synthetic caption-based supervision. Notably, RankCLIP shows the most significant improvement in top-1 accuracy, reinforcing its practical advantages where the highest-ranked prediction is most critical.
4.3 Zero-shot Cross-modal Retrieval
We further evaluate RankCLIP on zero-shot cross-modal retrieval tasks, including image-to-text and text-to-image retrieval, using the MSCOCO [30] dataset. As shown in Table 1, RankCLIP outperforms all baselines, though the improvements are less pronounced compared to zero-shot classification. This smaller margin of improvement may be attributed to the increased complexity of retrieval tasks, which require fine-grained image-text alignment across varying resolutions and object details – challenges distinct from the more direct pattern recognition in classification.
@@ 4.4 Robustness to Distribution Shifts
To evaluate the robustness of RankCLIP under distribution shifts, we test all approaches on three variants of ImageNetV2 [45] and ImageNet-R [22], which assess resilience to different real-world deviations from ImageNet1K. As shown in Table 2, RankCLIP consistently outperforms all baselines, achieving the highest accuracy across all datasets. These results indicate that RankCLIP not only improves standard zero-shot classification but also enhances adaptation to real-world distribution shifts.